OpenStudy (anonymous):

In centroid D is on median AD = x + 3 and DM = 2x - 1. Find AM. A. 7 B. 5/3 C. 8 D. 3 1/2

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

can you post a picture?

OpenStudy (mathmale):

Yes, I strongly suggest posting any illustration that came with this problem, or drawing your own illustration. Label it.

pooja195 (pooja195):

Do you mean this? 

OpenStudy (anonymous):

it didnt come with an illustration

OpenStudy (dayakar):

 AD: DM =2:1 \[\frac{ x+3 }{ 2x-1 }=\frac{ 2 }{ 1 }\] now find x